### v1

# Prodiabatic Elimination: Higher Order Elimination of Fast Variables with Quantum Noise

Abstract
We introduce the prodiabatic elimination, a powerful approximation technique that systematically extends the adiabatic elimination of fast degrees of freedom in light-matter coupled systems. Through a controlled expansion of operators, the prodiabatic elimination incorporates higher-order corrections and consistently includes noise contributions, leading to a significantly improved performance compared to standard adiabatic elimination. Importantly, it retains the simplicity and computational efficiency of the adiabatic elimination, making it convenient for practical applications. We demonstrate the approach on two setups: a driven dissipative Jaynes-Cummings model and a three-level system in a two-mode cavity that performs stimulated Raman adiabatic passage (STIRAP). These examples establish the prodiabatic elimination as a robust and broadly applicable tool for analyzing open quantum systems.
preprint: APS/123-QED
Introduction– Exact descriptions of open quantum dynamics are rarely available, making approximation methods essential for constructing effective models. Systems exhibiting well-separated time scales are particularly amenable to such approaches, as the separation can be exploited to derive simplified effective models [6], with applications in physics [34, 11, 2], chemistry [20], biology [8], medicine [22], and even finance [19]. One of the most prominent approaches to tackle systems with a separation of timescales is provided by the adiabatic elimination [17, 34, 28, 26, 11, 15, 38, 35], in which fast degrees of freedom are systematically removed, yielding an effective description for the remaining slow variables.
Manipulating quantum systems with light is a prominent approach both to investigate fundamental physics, and to implement quantum technologies. Fields relying on this paradigm include cavity [33] and circuit QED [9], optomechanics [4, 23], and cavity materials engineering [27, 21, 31]. It is often desired for the light to react fast, enabling, for instance, efficient readout of a quantum system [3]. In this case, the light can be adiabatically eliminated [34] leading to a much simpler description of the remaining quantum system, see Fig. 1 for an illustration of the setup. The adiabatic elimination is widely used because it is simple to apply and may yield analytical expressions with strong predictive power.
Typically, the adiabatic elimination is applied to leading order in the separation of timescales, which limits its applicability. Going beyond leading order has proven to be a challenging task. A central challenge thereby is the systematic treatment of noise, a problem well known in classical stochastic systems [36, 40, 18]. While different approaches exist [5, 29, 32, 37], they are significantly more difficult to employ than the leading order one, making them much less accessible. Importantly, these approaches focus on single-time observables, leaving the evaluation of multi-time correlation functions, such as the prominent $g^{(2)}$-function from quantum optics [39], an open problem.
Figure 1: General setup. A quantum system (purple) is embedded in a driven cavity. The quantum system is assumed to evolve slowly compared to the time-scale governing the cavity, allowing to eliminate the light. The cavity drive is denoted by $f(t)$, its linewidth by $\kappa$, and the quantum system dissipates at rate $\gamma$.
Here, we overcome these difficulties by introducing the prodiabatic elimination. In contrast to previous approaches that go beyond leading order, which rely on the master equation, the prodiabatic elimination is based on the quantum Langevin equation, which naturally allows for describing multi-time correlation function. Moreover, focusing on normal- and time-ordered observables in a zero-temperature environment makes the prodiabatic elimination comparably simple to employ, without losing experimental relevance. We stress that the prodiabatic elimination fully takes into account contributions from the vacuum noise, which we show are relevant in the short-time dynamics of multi-time observables.
The prodiabatic elimination– We model the dynamics of the cavity field with the quantum Langevin equation [16] ($\hbar=1$)
$\dot{\hat{a}}=-\left(\frac{\kappa}{2}+i\Delta\right)\hat{a}-ig\hat{b}+if(t)-\sqrt{\kappa}\hat{a}_{\rm in}\,$, (1)
where $\kappa$ is the cavity dissipation rate, $\Delta$ is the cavity frequency (or detuning in a rotating frame), $g$ denotes the coupling to the intracavity system which we generally refer to as atom, $f(t)$ is a (possibly time-dependent) coherent drive amplitude, and $\hat{a}_{\rm in}$ is an input operator of zero mean, describing the vacuum noise entering the cavity. The operator $\hat{b}$ acts on the atom and is assumed to be governed by the equation of motion
$\dot{\hat{b}}=-i\Omega\hat{v}-\hat{r}\left(\frac{\gamma}{2}\hat{b}+ig\hat{a}+\sqrt{\gamma}\hat{b}_{\rm in}\,\right),$ (2)
where $i\Omega\hat{v}$ describes the evolution of $\hat{b}$ in the absence of the cavity and dissipation, $\gamma$ denotes the atomic dissipation rate, $\hat{b}_{\rm in}$ is a zero mean input operator, describing vacuum noise, and $\hat{r}=[\hat{b},\hat{b}^{\dagger}]$. The terms proportional to $g$ in Eqs. (1, 2) result from a Jaynes-Cummings-like coupling $g(\hat{a}\hat{b}^{\dagger}+\hat{b}\hat{a}^{\dagger})$. A Markovian master equation that is equivalent to these Langevin equations is provided in the Supplemental Material [1].
We introduce a small parameter $\epsilon$ which ensures a separation of timescales between the cavity, evolving fast, and the atom, evolving slowly;
$\mathcal{O}\left(\frac{\gamma}{\kappa}\right)=\mathcal{O}\left(\frac{\Omega}{\kappa}\right)=\epsilon^{2}\,$, $\mathcal{O}\left(\frac{g}{\kappa}\right)=\mathcal{O}\left(\frac{f}{\kappa}\right)=\epsilon\,.$ (3)
The conditions, apart from $\Omega/\kappa$, are known as the bad-cavity limit [34], implying the Purcell factor $F_{\rm p}=\frac{4g^{2}}{\gamma\kappa}$ to be of order $1$. The last condition in Eq. (3) may be relaxed to $\mathcal{O}\left(f/\kappa\right)=1$, see [1].
Our goal is to evaluate time- and normal-ordered expectation values of the form
$\mathcal{G}(\{\tau_{m},t_{n}\})=\expectationvalue{\hat{a}^{\dagger}(\tau_{1})\cdots\hat{a}^{\dagger}(\tau_{M})\hat{\Sigma}(t)\hat{a}(t_{N})\cdots\hat{a}(t_{1})}\,$, (4)
where $\hat{\Sigma}(t)$ is an unspecified atom operator and time ordering means $t_{1}\leq\cdots\leq t_{N}$ as well as $\tau_{1}\leq\cdots\leq\tau_{M}$. Such expectation values are of central importance in quantum optics, as they encode measurable photon correlations including the first and second order coherence functions.
The leading-order adiabatic elimination can be obtained by setting $\dot{\hat{a}}=0$ in Eq. (1) and solving for $\hat{a}$. Dropping the noise term, which is justified for time- and normal-ordered expectation values [34], results in $\hat{a}=\hat{a}_{\rm adb}=-\frac{2it_{\rm c}}{\kappa}[g\hat{b}(t)-f(t)]$,
where $t_{\rm c}=\left(1+\frac{2i\Delta}{\kappa}\right)^{-1}$ is the cavity susceptibility. Note that $\hat{a}_{\rm adb}$ is an atomic operator, it does not act on the cavity. The correlation functions in Eq. (4) may then be evaluated using the quantum regression theorem [24, 10], see also Ref. [1].
To go beyond the adiabatic elimination we follow Ref. [34] and consider the formal solution of Eq. (1) (with initial condition in the far past)
$\hat{a}(t)=-ig\int_{0}^{\infty}d\tau\;e^{-\frac{\kappa\tau}{2t_{\rm c}}}\hat{b}(t-\tau)+iF(t)+\hat{A}_{\rm in}(t)\,$, (5)
where
$\hat{A}_{\rm in}(t)=$ $-\sqrt{\kappa}\int_{0}^{\infty}d\tau\;e^{-\frac{\kappa\tau}{2t_{\rm c}}}\hat{a}_{\rm in}(t-\tau)$ (6)
$F(t)=$ $\int_{0}^{\infty}d\tau\;e^{-\frac{\kappa\tau}{2t_{\rm c}}}f(t-\tau).$ (7)
For a time-independent drive, we find $F=\frac{2t_{\rm c}}{\kappa}f$. Due to the separation of time-scales, $\hat{b}$ evolves on a time-scale much slower than the damping governed by $1/\kappa$. We may therefore expand $\hat{b}(t-\tau)$ with respect to $\tau$ in Eq. (5) as outlined in the supplemental material [1]. We find that to evaluate the expectation value in Eq. (4) up to order $\epsilon^{N+M+2}$, we may approximate the cavity annihilation operator as
$\hat{a}(t)$ $\simeq\hat{a}_{\rm pdb}(t)+\hat{A}_{\rm in}(t)$ (8)
$-2\sqrt{\kappa}\hat{B}(t)\int_{0}^{\infty}d\tau\frac{e^{-\frac{\kappa\tau}{2t^{*}_{\rm c}}}-e^{-\frac{\kappa\tau}{2t_{\rm c}}}}{|t_{\rm c}|^{2}-t_{\rm c}^{2}}\hat{a}_{\rm in}^{\dagger}(t-\tau)\,$,
where the right-hand side only involves atom operators and cavity input operators. The substitution in Eq. (8) is at the heart of the prodiabatic elimination. The first term on the right-hand side is an operator that only acts on the atom
$\begin{split}\hat{a}_{\rm pdb}(t)=&\left[1+4\frac{g^{2}t_{\rm c}^{2}}{\kappa^{2}}\hat{r}(t)\right]\left[-\frac{2igt_{\rm c}}{\kappa}\hat{b}(t)+iF(t)\right]\\ &-2i\frac{\gamma gt_{\rm c}}{\kappa^{2}}\hat{r}(t)\hat{b}(t)+4\frac{\Omega gt_{\rm c}^{2}}{\kappa^{2}}\hat{v}(t)\,.\end{split}$ (9)
The leading order terms in this operator reduce to $\hat{a}_{\rm adb}$ for time-independent drives. However, in contrast to the adiabatic elimination, $\hat{a}_{\rm pdb}$ contains terms up to order $\epsilon^{3}$. These higher order terms not only modify the pre-factor of the $\hat{b}$ operator appearing already in $\hat{a}_{\rm adb}$, but they also introduce additional atom operators. While in the adiabatic elimination, photons emitted from the atom into the cavity quickly leave the cavity, this is no longer true in the prodiabatic elimination. As a consequence, removing a photon from the cavity is no longer equivalent to the action of the (displaced) atom annihilation operator $\hat{b}$, as predicted by $\hat{a}_{\rm adb}$.
The second and third term on the right-hand side of Eq. (8) are due to noise, where
$\hat{B}=4|t_{\rm c}|^{2}t_{\rm c}(2|t_{\rm c}|^{2}-t_{\rm c}^{2})\frac{g^{2}}{\kappa^{3}}$ (10)
[…]
is an atom operator that is order $\epsilon^{4}$. Note that the term $\hat{A}_{\rm in}$ ensures that $\hat{a},\hat{a}^{\dagger}]=1$. We now insert Eq. (8) into Eq. (4) and only keep terms up to order $\epsilon^{N+M+2}$. Using the commutator (for $t\geq t^{\prime}$) $\hat{A}_{\rm in}(t),\hat{a}(t^{\prime})]=e^{-\frac{(t-t^{\prime})\kappa}{2t_{\rm c}}}\hat{B}(t^{\prime})+\mathcal{O}\left(\epsilon^{5}\right)$, we then find [1]
$\mathcal{G}\left(\{\tau_{m},t_{n}\}\right)=\bigg\langle\hat{a}_{\rm pdb}^{\dagger}(\tau_{1})\cdots\hat{a}_{\rm pdb}^{\dagger}(\tau_{M})\hat{\Sigma}(t)\hat{a}_{\rm pdb}(t_{N})\cdots\hat{a}_{\rm pdb}(t_{1})\bigg\rangle$ (11)
$+\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}e^{-\frac{\kappa(t_{j}-t_{i})}{2t_{\rm c}}}\bigg\langle\mathcal{T}_{\mathcal{G}}\hat{a}_{\rm pdb}^{\dagger}(\tau_{1})\cdots\hat{a}_{\rm pdb}^{\dagger}(\tau_{M})\hat{\Sigma}(t)\hat{B}(t_{i})\prod_{n\neq i,j}^{N}\hat{a}_{\rm pdb}(t_{n})\bigg\rangle$
$+\sum_{i=1}^{M-1}\sum_{j=i+1}^{M}e^{-\frac{\kappa(\tau_{j}-\tau_{i})}{2t_{\rm c}^{*}}}\bigg\langle\mathcal{T}_{\mathcal{G}}\hat{B}^{\dagger}(\tau_{i})\bigg[\prod_{k\neq i,j}^{M}\hat{a}_{\rm pdb}^{\dagger}(\tau_{k})\bigg]\hat{\Sigma}(t)\hat{a}_{\rm pdb}(t_{N})\cdots\hat{a}_{\rm pdb}(t_{1})\bigg\rangle\,+\mathcal{O}(\epsilon^{N+M+3}),$
where $\mathcal{T}_{\mathcal{G}}$ ensures time-ordering according to Eq. (4), i.e., operators with $t_{n}$ are time-ordered and on the right of $\hat{\Sigma}(t)$ and operators with $\tau_{m}$ are anti-time ordered and to the left of $\hat{\Sigma}(t)$. We note that as long as we consider times that are well separated with respect to the fast time-scale, $\kappa|t_{i}-t_{j}|,\,\kappa|\tau_{l}-\tau_{m}|\gg 1$, we may neglect the noise contributions provided by the second and third lines in Eq. (11). It is thus often sufficient to only retain the first line, as illustrated in Fig. 2 (b) below.
Similar to the adiabatic elimination, the prodiabatic elimination thus expresses all correlation functions in terms of the atom alone. To evaluate Eq. (11), one may start by considering single-time expectation values. Using Eq. (2), one may express $\langle\dot{\hat{b}}\rangle$ in terms of expectation values that are all of the form of Eq. (4). These can then be expressed through atom operators alone using Eq. (11). In this way, one may find a closed set of equations for single-time expectation values that can be solved. Multi-time expectation values may then be evaluated using the quantum regression theorem [10, 25].
Figure 2: Prodiabatic elimination for the driven-dissipative Jaynes-Cummings model. (a) Time-dependent behavior of $\langle\hat{\sigma}_{z}\rangle$ for different drive strengths. Initially the atom is in the ground state. (b) Second order correlation function. Semi-transparent: numerical solution, solid: prodiabatic elimination, dashed adiabatic elimination. The inset in shows the short timescale behavior where the noise contribution is relevant. Parameters: $g/\kappa=3/20$, $\gamma/\kappa=5\cdot 10^{-3}$, $\Omega/\kappa=5\cdot 10^{-4}$, $\Delta/\kappa=1/20$, in (b) $f/\kappa=5/2\cdot 10^{-4}$.
Jaynes-Cummings model–As a first example, we consider a two-level system as the atom with $|\hat{b}=\hat{\sigma}\equiv\outerproduct{0}{1}$ and $\hat{v}=\hat{\sigma}$. Eq. (2) then reduces to
$\dot{\hat{\sigma}}=-\left(\frac{\gamma}{2}+i\Omega\right)\hat{\sigma}+ig\hat{\sigma}_{z}\hat{a}+\sqrt{\gamma}\hat{\sigma}_{z}\hat{b}_{\rm in}\,$, (12)
[…]
$\dot{\hat{\rho}}=-i[\hat{H},\hat{\rho}]+\kappa\mathcal{D}[\hat{a}]\hat{\rho}+\gamma\mathcal{D}[\hat{\sigma}]\hat{\rho}\,$, (13)
with the Hamiltonian
$\hat{H}=\Delta\hat{a}^{\dagger}\hat{a}+\frac{\Omega}{2}\hat{\sigma}_{z}+g\left(\hat{a}^{\dagger}\hat{\sigma}+\hat{\sigma}^{\dagger}\hat{a}\right)-f\left(\hat{a}+\hat{a}^{\dagger}\right)\,$, (14)
and $\mathcal{D}[\hat{A}]\hat{\rho}=\hat{A}\hat{\rho}\hat{A}^{\dagger}+\frac{1}{2}\{\hat{A}^{\dagger}\hat{A},\hat{\rho}\}$. The dynamics of the atom are fully characterized by the equations for the averages
$\langle\dot{\hat{\sigma}}\rangle$ $=-\left(\frac{\gamma}{2}+i\Omega\right)\expectationvalue{\hat{\sigma}}+ig\expectationvalue{\hat{\sigma}_{z}\hat{a}}\,$, (15)
[…]
where all averages are in the form of Eq. (4). We may thus employ Eq. (11), noting that only the first line contributes as the expectation values in Eqs. (15) and (16) only feature a single annihilation or creation operator. From Eq. (9), we find
$\hat{a}_{\rm pdb}=-\frac{2igt_{\rm c}}{\kappa}\left[\left(1+\frac{\gamma}{\kappa}\frac{t_{\rm c}}{t_{\rm q}}+4\frac{g^{2}t_{\rm c}^{2}}{\kappa^{2}}\right)\hat{\sigma}+4\frac{gft_{\rm c}^{2}}{\kappa^{2}}\hat{\sigma}_{z}-\frac{f}{g}\right]\,$, (17)
where $t_{\rm q}=\left(1+\frac{2i\Omega}{\gamma}\right)^{-1}$ is the susceptibility of the atom. Using Eq. (11), Eqs. (15) and (16) reduce to
$\langle\dot{\hat{\sigma}}\rangle$= $-\frac{\Gamma}{2}\expectationvalue{\hat{\sigma}}-\frac{2gft_{\rm c}}{\kappa}\left(\expectationvalue{\hat{\sigma}_{z}}-\frac{\gamma t_{\rm c}^{2}}{\kappa}F_{\rm p}\right)\,$, (18)
[…]
where $\Gamma=\frac{\gamma}{t_{q}}\,(1+t_{\rm c}t_{\rm q}F_{\rm p})\left(1+\frac{\gamma t_{\rm c}^{2}}{\kappa}F_{\rm p}\right)$ and $F_{\rm p}=\frac{4g^{2}}{\gamma\kappa}$. These equations of motion fully determine the state of the atom.
The above equations of motion cannot be immediately reproduced by a Markovian master equation (a similar behavior is found in Ref. [37]). However, a master equation that produces the same solution up to $\epsilon^{3}$, i.e., the order to which the prodiabatic elimination can be trusted, can be found [1]; at resonance $\Delta=\Omega=0$, it reads
$\dot{\hat{\rho}}$= $-i[\hat{H}_{\rm pdb},\hat{\rho}]+\gamma(1+F_{\rm p})\left(\frac{\gamma}{\kappa}F_{\rm p}\mathcal{D}[\hat{\sigma}]+\mathcal{D}[\hat{\sigma}+\xi\hat{\sigma}_{z}]\right)\hat{\rho}\,$, (20)
with
$\hat{H}_{\rm pdb}=-\frac{2gf}{\kappa}\left(\frac{\gamma}{2\kappa}F_{\rm p}+1\right)\hat{\sigma}_{y},\hskip 9.95863pt\xi=\frac{2fg}{\kappa^{2}}\frac{F_{\rm p}}{F_{\rm p}+1},$ (21)
and $\hat{\sigma}_{y}=i(\hat{\sigma}-\hat{\sigma}^{\dagger})$. In the adiabatic elimination, where we ignore all terms that are proportional to $\gamma^{2}$ or $\gamma\xi$, the effect of the cavity on the atom is a coherent drive, as well as an enhanced decay $\gamma\rightarrow\gamma(1+F_{\rm p})$ [34]. In the prodiabatic elimination, on the other hand, we find the same effects but with modified parameters. Additionally we find dissipation with the jump operator $\hat{\sigma}+\xi\hat{\sigma}_{z}$. This jump does not result in perfect relaxation, but instead takes the atom to a point on the Bloch sphere that is close to the ground state. We interpret this jump as arising from the possibility of re-absorbing a photon that is emitted from the atom into the cavity. In contrast to the adiabatic elimination, the prodiabatic elimination takes this process into account perturbatively. As illustrated in Fig. 2 (a) the prodiabatic elimination captures the time-dependent behavior of $\langle\hat{\sigma}_{z}\rangle$ well for small drives and starts to break down as $f$ increases. Compared to the adiabatic elimination, the prodiabatic elimination remains valid to larger drive strengths.
Using the quantum regression theorem together with the prodiabatic elimination, we found an analytical expression for the second order coherence function
$g^{(2)}(t)=\frac{\langle\hat{a}^{\dagger}(0)\hat{a}^{\dagger}(t)\hat{a}(t)\hat{a}(0)\rangle}{\langle\hat{a}^{\dagger}(0)\hat{a}(0)\rangle^{2}}\,.$ (22)
As illustrated in Fig. 2 (b) the prodiabatic elimination captures the time-dependent behavior of the $g^{(2)}$-function more adequately compared to the adiabatic elimination. In the low drive limit $f^{2}/(\kappa\gamma)\ll 1$ as well as on resonance ($\Delta=\Omega=0$), we find
$\begin{split}g^{(2)}_{\rm pdb}(t)=&\left[1-F_{\rm p}^{2}\left(1-\frac{\gamma^{2}}{\kappa^{2}}\left(F_{\rm p}+1\right)^{2}\right)e^{-\frac{\Gamma t}{2}}\right]^{2}\\ &+2e^{-\frac{\kappa t}{2}}\frac{\gamma}{\kappa}F_{\rm p}^{2}\left(1+F_{\rm p}\right)\left(1-F_{\rm p}^{2}e^{-\frac{\Gamma t}{2}}\right).\end{split}$ (23)
The result of the adiabatic elimination is recovered by dropping all terms proportional to $\gamma/\kappa$. The second line of Eq. (23) corresponds to the short-time correction arising from the noise terms, i.e., the second and third line of Eq. (11).
[…]
In this case, the atom is described by transition (and population) operators $\hat{\sigma}_{ij}=|i\rangle\langle j|$, with $i,j=1,2,3$, and the cavity modes have associated annihilation operators $\hat{a}_{\rm H}$, $\hat{a}_{\rm v}$. Equations (1) and (2) still describe this scenario if we consider the constituents to be vectors, i.e., $\hat{b}=(\hat{\sigma}_{13},\hat{\sigma}_{23})$, $\hat{a}=(\hat{a}_{\rm H},\hat{a}_{\rm V})$, and similarly for $\hat{a}_{\rm in}$ and $\hat{b}_{\rm in}$. Additionally, $\hat{r}$ becomes a two by two matrix such that $\hat{r}_{l,k}=[\hat{b}_{l},\hat{b}^{\dagger}_{k}]$. For simplicity, we consider $\Omega=\Delta=0$ in this section. In this case, Eqs. (1) and (2) are then equivalent to the master equation [1]
$\dot{\hat{\rho}}=-i[\hat{H},\hat{\rho}]+\kappa\sum_{\mu\in\{H,V\}}\mathcal{D}[\hat{a}_{\mu}]\hat{\rho}+\gamma\sum_{j\in\{1,2\}}\mathcal{D}[\hat{\sigma}_{j3}]\hat{\rho}$ (24)
where $\kappa$ denotes the linewdith of the cavity (assumed equal for the two modes), $\gamma$ the dissipation strength of the three-level system (assumed equal for the two transitions) and the Hamiltonian reads
$\hat{H}=g(\hat{a}_{\rm V}^{\dagger}\hat{\sigma}_{23}+\hat{a}_{\rm H}^{\dagger}\hat{\sigma}_{13})-f_{\rm V}(t)\hat{a}_{\rm V}^{\dagger}-f_{\rm H}(t)\hat{a}_{\rm H}^{\dagger}+\text{h.c.}\,$, (25)
with $f_{\rm H}(t)$ and $f_{\rm V}(t)$ being time-dependent drive envelope functions, allowing to apply pulses to the H and V-mode respectively.
By prodiabatically eliminating the cavity modes, we find the (unnormalized) dark state
[…]
Refer to captionRefer to caption
Figure 3: Prodiabatic elimination for STIRAP protocol. Semi-transparent: numerical simulations, dashed: adiabatic elimination, solid: prodiabatic elimination and doted: Lindblad master equation Eq. (S34). The drives are boxcar functions: $f_{\rm V}(t)=\kappa$ if $|55-\kappa t|\leq 10$ and $f_{\rm H}(t)=\kappa$ if $|45-\kappa t|\leq 10$, further $g/\kappa=1/10$ and $\gamma/\kappa=5\cdot 10^{-4}$.
Conclusions and outlook.—The prodiabatic elimination provides analytical expressions capturing a wealth of phenomena beyond the standard adiabatic elimination. This includes computing correlation functions beyond leading order, capturing short-time behavior via vacuum noise contributions, as well as the combined effect of pulse modulation and cavity filtering.
These capabilities pave the way for the development of optimal control protocols in cavity QED [12, 14, 13, 30], which strongly rely on the interplay between shaped pulses and cavity filtering. Future extensions of the method may include treating thermal noise, as well as a general procedure to obtain Lindblad master equations.
Supplemental Material:
Prodiabatic Elimination:
Higher Order Elimination of Fast Variables with Quantum Noise
Jan Neuser1,2, Marcelo Janovitch1, Matteo Brunelli3, and Patrick P. Potts1
1Department of Physics and Swiss Nanoscience Institute,
University of Basel, Klingelbergstrasse 82, 4056 Basel, Switzerland
2 Atominstitut, TU Wien, 1020 Vienna, Austria
3JEIP, UAR 3573 CNRS, Collège de France, PSL Research University,
11 Place Marcelin Berthelot, 75321 Paris Cedex 05, France
In this supplementary material, we give additional formulas and derivations supporting the statements of the main text including: the general form of the Lindblad master equation our approach is applicable to, the derivation of Eq. (11), the use of the quantum regression theorem for multi-time correlation functions, the Lindblad master equations equivalent to the dynamics of the atom after the prodiabatic elimination for both examples considered, a further exploration of parameter space for STIRAP, and extending the approach to stronger drives.
I 1. Equivalence between Lindblad Master equation and Langevin equations
Equations (12) can be obtained from the general quantum Langevin equation in the presence of two input channels,
[…]
$-[\hat{\mathcal{O}},\hat{b}^{\dagger}]\quantity(\frac{\gamma}{2}\hat{b}+\sqrt{\gamma}\hat{b}_{\text{in}})-\quantity(\frac{\gamma}{2}b^{\dagger}+\sqrt{\gamma}\hat{b}_{\text{in}}^{\dagger})[\hat{\mathcal{O}},\hat{b}$].,
where, the Hamiltonian is
$\hat{H}=\hat{H}_{\rm b}+\Delta\hat{a}^{\dagger}\hat{a}+g\left(\hat{a}\hat{b}^{\dagger}+\hat{b}\hat{a}^{\dagger}\right)-f(t)\left(\hat{a}+\hat{a}^{\dagger}\right)\,$, (S2)
and $\hat{H}_{\rm b}$ models the the atom in the absence of the cavity. This is done by setting the generic operator $\hat{\mathcal{O}}=\hat{a},~\hat{b}$, and introducing $\Omega\hat{v}=[\hat{b},\hat{H}_{\rm b}]$.
In turn, Eq. (S1) is equivalent to the master equation [16]
$\dot{\hat{\rho}}=-i[\hat{H},\hat{\rho}]+\kappa\mathcal{D}[\hat{a}]\hat{\rho}+\gamma\mathcal{D}[\hat{b}]\hat{\rho}\,$. (S3)
II 2. Derivation of the prodiabatic Elimination
The first step in deriving the prodiabatic elimination is to expand the the operator $\hat{b}(t-\tau)$ in the formal solution for $\hat{a}(t)$, see Eq. (5). To this end, we write
$\hat{b}(t-\tau)=\hat{b}(t)-\int_{0}^{\tau}dt^{\prime}\dot{\hat{b}}(t-t^{\prime})\,.$ (S4)
We now rewrite Eq. (2) as
$\dot{\hat{b}}(t)=-\kappa\hat{X}(t)-\hat{r}(t)\sqrt{\gamma}\hat{b}_{\rm in}(t),$ (S5)
with
$\hat{X}(t)=\frac{i\Omega}{\kappa}\hat{v}(t)+\frac{\gamma}{2\kappa}\hat{r}(t)\hat{b}(t)+\frac{ig}{\kappa}\hat{r}(t)\hat{a}(t)\,.$ (S6)
Inserting Eq. (S5) into Eq. (S4) allows us to write
$\hat{b}(t-\tau)=\hat{b}(t)+\kappa\tau\hat{X}(t)+\frac{\sqrt{\gamma}}{\kappa}\int_{0}^{\tau}\kappa dt^{\prime}\;\hat{r}(t-t^{\prime})\hat{b}_{\rm in}(t-t^{\prime})-\kappa\int_{0}^{\tau}dt^{\prime}\;\int_{0}^{t^{\prime}}dt^{\prime\prime}\dot{\hat{X}}(t-t^{\prime\prime})\,.$ (S7)
To make progress, we note that $\hat{b}=\mathcal{O}(1)$ and we anticipate that $\hat{a}=\mathcal{O}(\epsilon)$ [see Eq. (S9)]. As a consequence, $\hat{X}=\mathcal{O}(\epsilon^{2})$. We further note that $\kappa\tau$ and $\kappa dt$ are $\mathcal{O}(1)$ due to the exponentials appearing in the integrals of Eq. (5) and (S7). Since the atom evolves on the timescale of $\epsilon$ (or slower) compared to the cavity, we have $\dot{X}/\kappa=\mathcal{O}(\epsilon^{3})$.
II.1 2.1 Ignoring the noise
We first drop all contributions from noise terms $\hat{a}_{\rm in}$ and $\hat{b}_{\rm in}$. In the next section, we discuss how to properly include them. Ignoring the noise terms allows us to write
$\hat{b}(t-\tau)=\hat{b}(t)+\kappa\tau\hat{X}(t)+\mathcal{O}(\epsilon^{3}).$ (S8)
To lowest order in $\epsilon$, we may also drop the second term on the right-hand side. Inserting $\hat{b}(t-\tau)=\hat{b}(t)$ into Eq. (5) results in
$\hat{a}=-\frac{2igt_{\rm c}}{\kappa}\hat{b}(t)+iF(t)+\mathcal{O}(\epsilon^{3}).$ (S9)
To include higher order terms, we insert Eq. (S9) into the right-hand side of Eq. (S6). Inserting the resulting $\mathcal{O}(\epsilon^{2})$ expression for $\hat{X}$ into Eq. (S8), and inserting that expression into Eq. (5) then results in $\hat{a}(t)=\hat{a}_{\rm pdb}(t)+\mathcal{O}(\epsilon^{4})$, with $\hat{a}_{\rm pdb}$ given in Eq. (9) in the main text.
II.2 2.2 Including noise
We now consider the terms that arise due to vacuum noise, starting with the third term on the right-hand side of Eq. (S7). Inserting this term into Eq. (5) shows that this term contributes to the annihilation operator $\hat{a}$ as
$\hat{\xi}(t)=-2\sqrt{\gamma}t_{\rm c}\left(\frac{g}{\kappa}\right)^{2}\int\kappa d\tau e^{-\frac{\kappa\tau}{2t_{\rm c}}}\int_{0}^{\tau}\kappa dt^{\prime}\hat{r}(t-t^{\prime})\hat{b}_{\rm in}(t-t^{\prime}).$ (S10)
This operator then contributes to the expectation values of interest as $\langle\cdots\hat{\xi}(t)\hat{a}(t_{k})\cdots\hat{a}(t_{1})\rangle$. We now show that this contribution can be neglected. To this end, we consider
$\hat{b}_{\rm in}(t-t^{\prime}),\hat{a}(t_{k})]=-\frac{2igt_{\rm c}}{\kappa}[\hat{b}_{\rm in}(t-t^{\prime}),\hat{b}(t_{k})]+\mathcal{O}(\epsilon^{3}),$ (S11)
where we used Eq. (S9). This commutator is only finite when $t-t^{\prime}\leq t_{k}$, as system operators can only be influenced by input operators from the past [16]. Using Eq. (S7), we may write
$\hat{b}(t_{k})=\hat{b}(t-t^{\prime}-\delta)-\kappa(t_{k}-t+t^{\prime}+\delta)\hat{X}(t-t^{\prime}-\delta)-\frac{\sqrt{\gamma}}{\kappa}\int_{t-t^{\prime}-\delta}^{t_{k}}\kappa d\tau\hat{r}(\tau)\hat{b}_{\rm in}(\tau)+\mathcal{O}(\epsilon^{3}),$ (S12)
where $\delta$ is an arbitrary small time. Inserting this expression into Eq. (S11), we find
$\hat{b}_{\rm in}(t-t^{\prime}),\hat{a}(t_{k})]=\mathcal{O}(\epsilon^{3}).$ (S13)
As a consequence, we may commute the input noise operators appearing in $\langle\cdots\hat{\xi}(t)\hat{a}(t_{k})\cdots\hat{a}(t_{1})\rangle$ all the way to the right, where they annihilate the vacuum. The correction to Eq. (4) arising from the $\epsilon^{3}$ term in Eq. (S13) contributes to the order $\epsilon^{N+M+3}$ and can be neglected as our approximation is only valid to order $\epsilon^{N+M+2}$.
Next we consider the effect of the vacuum input noise of the cavity, which does contribute to Eq. (4) through the second and third line of Eq. (11). From Eq. (5), we can infer the contributions from the cavity noise to enter through terms of the form $\langle\cdots\hat{A}_{\rm in}(t)\hat{a}(t_{k})\cdots\hat{a}(t_{1})\rangle$. We therefore consider
$\hat{A}_{\rm in}(t),\hat{a}(t_{k})]=i\frac{g}{\kappa}\int_{0}^{\infty}\kappa\tau\int_{0}^{\infty}\kappa\tau^{\prime}e^{-\frac{\kappa}{2t_{\rm c}}(\tau+\tau^{\prime})}\frac{1}{\sqrt{\kappa}}[\hat{a}_{\rm in}(t-\tau),b(t_{k}-\tau^{\prime})],$ (S14)
where we used Eq. (5). Note that the right-hand side is only finite for $t-\tau\leq t_{k}-\tau^{\prime}$. To make further progress, we rewrite Eq. (S7) as
$\hat{b}(t^{\prime})=\hat{b}(t-\delta)-\kappa(t^{\prime}-t+\delta)\hat{X}(t-\delta)-\int_{t-\delta}^{t^{\prime}}\kappa d\tau\;\int_{t-\delta}^{\tau}\kappa d\tau^{\prime}\dot{\hat{X}}(\tau^{\prime})/\kappa\,$, (S15)
where this time we dropped the noise term proportional to $\hat{b}_{\rm in}$ as it will not contribute. Equation (S15) allows us to write
$\hat{a}_{\rm in}(t),\hat{b}(t^{\prime})]=-\int_{t-\delta}^{t^{\prime}}\kappa d\tau\int_{t-\delta}^{\tau}\kappa d\tau^{\prime}[\hat{a}_{\rm in}(t),\dot{\hat{X}}(\tau^{\prime})/\kappa].$ (S16)
To evaluate this expression, we need
$\dot{\hat{X}}(t)=\frac{g}{\kappa}\hat{a}^{\dagger}(t)\left[[\hat{r}(t),\hat{b}(t)]\left(g\hat{a}(t)-i\frac{\gamma}{2}\hat{b}(t)\right)+\Omega[\hat{v}(t),\hat{b}(t)]\right]+\cdots,$ (S17)
where we only wrote out the terms that contribute to Eq. (S16). Using
$\hat{a}^{\dagger}(t)=-\sqrt{\kappa}\int_{0}^{\infty}d\tau e^{-\frac{\kappa\tau}{2t_{\rm c}^{*}}}\hat{a}_{\rm in}^{\dagger}(t-\tau)+\mathcal{O}(\epsilon),$ (S18)
we then find
$\frac{1}{\sqrt{\kappa}}[\hat{a}_{\rm in}(t),\dot{\hat{X}}(\tau^{\prime})/\kappa]=-\frac{g}{\kappa^{2}}e^{-\kappa\frac{\tau^{\prime}-t}{2t_{\rm c}^{*}}}\left[[\hat{r}(t),\hat{b}(t)]\left(g\hat{a}(t)-i\frac{\gamma}{2}\hat{b}(t)\right)+\Omega[\hat{v}(t),\hat{b}(t)]\right]+\mathcal{O}(\epsilon^{4}),$ (S19)
for $t\leq\tau^{\prime}$ and zero otherwise. Inserting Eq. (S19) into Eq. (S16), and inserting that equation into Eq. (S14) we find
$\hat{A}_{\rm in}(t),\hat{a}(t^{\prime})]=e^{-\frac{(t-t^{\prime})\kappa}{2t_{\rm c}}}\hat{B}(t^{\prime})+\mathcal{O}\left(\epsilon^{5}\right),$ (S20)
for $t\geq t^{\prime}$, where $\hat{B}(t)$ is given in Eq. (10) in the main text.
Knowing how the noise terms contribute to the correlation functions in Eq. (4), we can find the appropriate expansion of $\hat{a}$ by approximating Eq. (S7) with Eqs. (S17) and (S18), dropping terms including $\hat{b}_{\rm in}$, and inserting back into the formal solution in Eq. (5). This results in Eq. (8) in the main text, which respects Eq. (S20). To find the correlation functions in Eq. (4) up to order $\epsilon^{N+M+2}$, we insert the expansion of $\hat{a}$ in Eq. (8) into Eq. (4) and use the commutation relation in Eq. (S20) repeatedly, only keeping terms up to the desired order. This results in our main result, Eq. (11) in the main text.
III 2. Lindblad master equations equivalent to the prodiabatic elimination
Here we give the full expression of the master equation for the two models considered in the main text.
III.1 2.1 Jaynes-Cummings model
The equations of motion found for the atom through the prodiabatic elimination, see Eqs. (18) and (19), can be reproduced by a Lindblad master equation up to the relevant order. This equation reads
$\dot{\hat{\rho}}=-i[\hat{H}_{\text{pdb}},\hat{\rho}]+\Gamma_{1}\mathcal{D}[\hat{\sigma}]\hat{\rho}+\Gamma_{0}D[\hat{\sigma}+\xi\hat{\sigma}_{z}]\hat{\rho}\,$, (S21)
where
$\hat{H}_{\rm pdb}=c_{\rm z}\hat{\sigma}_{\rm z}+c_{\rm x}\hat{\sigma}_{\rm x}+c_{\rm y}\hat{\sigma}_{\rm y}$ (S22)
with $\sigma_{z}=[\sigma^{\dagger},\sigma]$, $\sigma_{\rm x}=\sigma+\sigma^{\dagger}$ and $\hat{\sigma}_{y}=i(\hat{\sigma}-\hat{\sigma}^{\dagger})$. Further, the coefficients read
[…]
The rates in Eq. (S21) are given by
[…]
Equation (S21) will reproduce the equations of motion in Eqs. (18) and (19), up to terms that are proportional to $\gamma\epsilon^{4}$, which can safely be neglected. We note that this term is however necessary to obtain Lindblad form, which ensures complete positivity for any choice of parameters. Setting $\Delta=\Omega=0$, we recover the master equation in Eq. (20) in the main text.
III.2 2.2 STIRAP
For the STIRAP setup, the prodiabatic elimination results in the following equations of motion for the atom
$\langle\dot{\sigma}_{11}\rangle$ $=\Gamma_{\rm s}\langle\sigma_{33}\rangle-gF_{\rm H}\left(1+\frac{\gamma}{\kappa}F_{\rm p}\right)\left(\langle\sigma_{13}\rangle+\langle\sigma_{31}\rangle\right)-\frac{4F_{\rm V}g^{3}}{\kappa^{2}}\left(\langle\sigma_{23}\rangle+\langle\sigma_{32}\rangle\right)$ (S29)
$\langle\dot{\sigma}_{22}\rangle$ $=\Gamma_{\rm s}\langle\sigma_{33}\rangle-\frac{4F_{\rm H}g^{3}}{\kappa^{2}}\left(\langle\sigma_{13}\rangle+\langle\sigma_{31}\rangle\right)-gF_{\rm V}\left(1+\frac{\gamma}{\kappa}F_{\rm p}\right)\left(\langle\sigma_{23}\rangle+\langle\sigma_{32}\rangle\right)$ (S30)
$\langle\dot{\sigma}_{12}\rangle$ $=-g\left(F_{\rm V}\sigma_{13}+F_{\rm H}\sigma_{32}\right)$ (S31)
$\langle\dot{\sigma}_{13}\rangle$ $=-\Gamma_{\rm s}\sigma_{13}+gF_{\rm H}\left(1+\frac{2\gamma}{\kappa}F_{\rm p}\right)\sigma_{11}+gF_{\rm V}\left(1+\frac{2\gamma}{\kappa}F_{\rm p}\right)\sigma_{12}-gF_{\rm H}\left(1-\frac{\gamma}{\kappa}F_{\rm p}\right)\sigma_{33}$ (S32)
$\langle\dot{\sigma}_{23}\rangle$ $=-\Gamma_{\rm s}\sigma_{23}+gF_{\rm V}\left(1+\frac{2\gamma}{\kappa}F_{\rm p}\right)\sigma_{22}+gF_{\rm H}\left(1+\frac{2\gamma}{\kappa}F_{\rm p}\right)\sigma_{21}-gF_{\rm V}\left(1-\frac{\gamma}{\kappa}F_{\rm p}\right)\sigma_{33}$ (S33)
where we introduced $\Gamma_{\rm s}=\gamma(1+F_{\rm p})(1+\frac{2\gamma}{\kappa}F_{\rm p})$ these equations are enough to define the equation of motion of a full basis, as we can use $\sigma_{ij}^{\dagger}=\sigma_{ji}$ as well as $\langle\sigma_{11}\rangle+\langle\sigma_{22}\rangle+\langle\sigma_{33}\rangle=1$. The adiabatic result can be obtained by removing all terms proportional to $\gamma/\kappa$ as well as doing the replacement $F_{i}\rightarrow 2f_{i}/\kappa$.
Up to terms of order $\gamma\epsilon^{4}$, these equations are reproduced by the master equation
[…]
with the Hamiltonian
$\begin{split}\hat{H}_{\rm pdb}=&-igF_{\rm H}\left(1+\frac{\gamma}{\kappa}F_{\rm p}\right)(\hat{\sigma}_{13}-\hat{\sigma}_{31})-igF_{\rm V}\left(1+\frac{\gamma}{\kappa}F_{\rm p}\right)(\hat{\sigma}_{23}-\hat{\sigma}_{32}).\end{split}$ (S35)
IV 3. STIRAP protocols with Gaussian pulses
In this section, we consider STIRAP protocols with different Gaussian pulses $f_{i}(t)=c_{i}\exp(-\frac{(t-\tau_{i})^{2}}{2s_{i}^{2}})$. The height and the width of the pulses, parametrized by $c_{i}$ and $s_{i}$, determine how fast the drive changes in time. To achieve ideal STIRAP, where the excited state is not populated, the drive needs to change sufficiently slowly. In the top of Fig. S1 we illustrate a close-to-ideal STIRAP protocol, whereas the bottom row illustrates that decreasing the width of the Gaussian pulses leads to a significant population of the third level.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure S1: Prodiabatic elimination for the STIRAP setup. Left panels: populations as a function of time. Semi-transparent: numerical simulations, dashed: adiabatic elimination and solid: prodiabatic elimination. The drives are Gaussian $f_{i}(t)=c_{i}\exp(-\frac{(t-\tau_{i})^{2}}{2s_{i}^{2}})$, where for the top: $c_{\rm H}/\kappa=c_{\rm V}/\kappa=3/4$, $\kappa s_{\rm H}=\kappa s_{\rm H}=12$, $\kappa\tau_{\rm V}=58$ and $\kappa\tau_{\rm H}=42$. For the bottom: $c_{\rm H}/\kappa=c_{\rm V}/\kappa=3/4$, $\kappa s_{\rm H}=\kappa s_{\rm H}=4$, $\kappa\tau_{\rm V}=52.5$ and $\kappa\tau_{\rm H}=47.5$. Right panels the corresponding pulses, solid $2f_{i}/\kappa$ (as they enter the adiabatic elimination) and dashed $F_{i}$ (as they enter the prodiabatic elimination). System parameters: $g/\kappa=1/5$ and $\gamma/\kappa=5\cdot 10^{-4}$